OpenStudy (anonymous):

The Chemistry departmental office has 800 black dry-erase markers. The 25 members of the department may take as many as they like. If 46 markers are taken the first day, how many different distributions to the faculty members are possible?

OpenStudy (anonymous):

@mukushla @eliassaab

OpenStudy (anonymous):

@siddhantsharan @hartnn

OpenStudy (anonymous):

Umm. Total markers chosen = 46 Let total markers chosen by person 1 = x1 -- (1 is in the subscript) by person 2 = x2 person 3 = x3..... . , person 25 = x25 Therefore x1 + x2 + x3 ........... + x25 = 46 Now, Find the non negative integral solutions for this using the formula.

OpenStudy (anonymous):

So it will be \[\left(\begin{matrix}46+25-1 \\ 25-1\end{matrix}\right)\]

OpenStudy (anonymous):

\[\left(\begin{matrix}70 \\ 24\end{matrix}\right)\]

OpenStudy (anonymous):

Can...we Do.this by..Permutation and combination.....Since..i have nt studied those..things which u posted nw

OpenStudy (anonymous):

LEmme try.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

Okay I can figure out 2 more ways to do this. First. Lets take a parallel problem to simplify matters.

OpenStudy (anonymous):

Let there be 3 members in the department and say 5 markers. Now these are 5 similar objects to be divided into three people. So lets consider two dividers. Basically, wherever we place this "divider" the objects are divided accordingly. For example: